Factor the polynomial completely ax – ay + bx - by
Solve 2x2 + 5x – 12 = 0

Answer by longjonsilver(2297)

1. a(x-y) + b(x-y)
this becomes (a-b)(x-y)
2. (2x-3)(x+4) = 0
so 2x-3=0 or x+4=0
so 2x=3 or x=-4
--> x=2/3, x=-4
jon.